# Training Set Reconstruction from Differentially Private Forests: How Effective is DP?

### v2

@@ Appendix E Additional Experiments on Scalability
In this section, we investigate the effect of the DP RFs training set size on the effectiveness of our dataset reconstruction attack. More precisely, we focus on DP RFs built of $\lvert\mathcal{T}\rvert=10$ trees of depth $5$, with DP budget $\varepsilon=5$. This configuration is representative of the experimental conditions presented in Section 4.1. We train such DP RFs with varying numbers of training examples $N\in\{25,50,100,200,300,400,500\}$. The remaining of the setup is as described in Section 4.1 (in particular, all results are averaged across $5$ different random seeds), apart from the maximum CPU time which is increased to $5$ hours.
@@
Table 5 displays the predictive accuracy of the built DP RFs, along with the success of the associated reconstruction attacks. Note that the results for the configuration used in Section 4.2 (i.e., $N=100$) are slightly better here because the CP formulations are often not solved to optimality and the solver benefits the additional running time to find slightly better solutions. This suggests that the results presented throughout the paper are likely to improve with increased adversarial computational power and advancements in CP solver technologies. We observe that as the training set size $N$ increases, the test set accuracy of the built DP RFs improves significantly. Meanwhile, the training accuracy decreases and converges toward the test accuracy, indicating better generalization of the built DP RFs. Another interesting phenomenon can be observed: the DP RFs training and test set accuracies sometimes increase simultaneously while increasing the training set size $N$ (e.g., from $N=300$ to $N=400$ for the Default Credit dataset). This is due to the fact that the relative impact of the added noise decreases when $N$ increases, allowing for better predictive performances. Crucially, this reduction in the relative effect of noise over the trees’ parameters also benefits the reconstruction attack: the reconstruction error significantly decreases as the size of the DP RF’s training set $N$ increases. For instance, while the reconstruction error from DP RFs trained on $N=100$ examples from the UCI Adult Income dataset is $12.3$% on average, this number decreases to $8.8$% (hence a relative decrease of $28$%) when it is trained on $N=500$ examples, although the number of examples to retrieve is significantly greater.
@@
Table 5: Results of our experiments on the reconstruction attack’s scalability. For varying training set sizes $N$, we report the predictive performances of the built DP RFs, along with the resulting reconstruction error (averaged over $5$ different random seeds).
@@
25 0.229 0.017 0.920 0.772
50 0.182 0.015 0.876 0.780
100 0.158 0.016 0.856 0.782
200 0.139 0.012 0.828 0.782
300 0.125 0.007 0.827 0.782
400 0.129 0.011 0.828 0.786
500 0.114 0.007 0.814 0.787
@@ Appendix F Additional Experiments on Partial Reconstruction
We report in Figure 9 the reconstruction error (measured for the unknown attributes) as a function of the number of known attributes. The results for all three datasets consistently show that the knowledge of a number of attributes does not help reconstruct the others. The slight increase in reconstruction error when the number of known attributes increases can be explained by the alignment process: as explained in Section 4.1, to figure out which reconstructed example corresponds to which original one, we first perform a minimum cost matching, before computing the average reconstruction error for each pair of original-reconstructed examples. In this partial reconstruction setup, this matching is still performed on the complete examples’s features vector even though the attack only retrieves part of their attributes. Hence, the improvement in reconstruction error induced by the matching process reduces as the number of unknown attributes decreases.
This observation empirically verifies the theoretical guarantees offered by DP: the protection enforced for the unknown attributes can not be “undone” through postprocessing the DP RF, even with the auxiliary information provided by the known attributes. In other words, DP prevents so-called linkage attacks [Dwork and Roth, 2014]. It is worth noting that these results contrast with the ones observed against non-DP RFs, for which knowledge of some of the attributes greatly helped reconstruct the others [Ferry et al., 2024b].
@@
Appendix G Broader Impact Statement
Machine learning models are increasingly trained on sensitive private data, such as medical records or personal histories. The need to audit their decision-making processes (e.g., to prevent discriminatory practices) has led to regulations advocating for greater transparency. To enable the safe release of statistics, model predictions or trained models, DP has become a widely adopted standard due to its appealing theoretical properties.
However, DP does not (at least directly) provide theoretical protection against dataset reconstruction attacks. Our work demonstrates that, in practice, machine learning models can comply with rigorous DP guarantees while still leaking a significant portion of their training data. This critical finding highlights the need for developing new defense mechanisms, and emphasizes the importance of the design of the DP mechanism itself, beyond the sole privacy budget. Going in this direction, our study provides insights into the impact of different DP design choices and advocates for empirically evaluating a model’s resilience against a variety of reconstruction attacks —including the one introduced in this work— before disclosure. Our findings emphasize that achieving privacy-preserving machine learning requires a holistic approach that carefully balances transparency, model utility, and empirical robustness against privacy threats. To this end, the proposed reconstruction attack provides a useful mathematically-rigorous way of benchmarking privacy protections.

### v4

@@ Appendix E Additional Experiments on Scalability
In this section, we investigate the effect of the DP RFs training set size on the effectiveness of our dataset reconstruction attack. More precisely, we focus on DP RFs built of $\lvert\mathcal{T}\rvert=10$ trees of depth $5$, with DP budget $\varepsilon=5$. This configuration is representative of the experimental conditions presented in Section IV-A. We train such DP RFs with varying numbers of training examples $N\in\{25,50,100,200,300,400,500\}$. The remaining of the setup is as described in Section IV-A (in particular, all results are averaged across $5$ different random seeds), apart from the maximum CPU time and the number of threads allocated to the solver, which are increased to $10$ hours and $32$ threads, respectively. This setup allows a thorough characterization of the solver anytime behaviour, even for the largest problems, as empirically evidenced in our results.
@@
Table VII displays the predictive accuracy of the built DP RFs, along with the success of the associated reconstruction attacks. Note that the results for the configuration used in Section IV-B (i.e., $N=100$) are slightly better here because the CP formulations are often not solved to optimality and the solver benefits the additional running time to find slightly better solutions. This suggests that the results presented throughout the paper are likely to improve with increased adversarial computational power and advancements in CP solver technologies. We observe that as the training set size $N$ increases, the test set accuracy of the built DP RFs improves significantly. Meanwhile, the training accuracy decreases and converges toward the test accuracy, indicating better generalization of the built DP RFs. Another interesting phenomenon can be observed: the DP RFs training and test set accuracies sometimes increase simultaneously while increasing the training set size $N$ (e.g., from $N=300$ to $N=400$ for the Default Credit dataset). This is due to the fact that the relative impact of the added noise decreases when $N$ increases, allowing for better predictive performances.
Crucially, this reduction in the relative effect of noise over the trees’ parameters also benefits the reconstruction attack: the reconstruction error significantly decreases as the size of the DP RF’s training set $N$ increases. For instance, while the reconstruction error from DP RFs trained on $N=100$ examples from the UCI Adult Income dataset is 12.312.3% on average, this number decreases to 8.58.5% (hence a relative decrease of 3131%) when it is trained on $N=500$ examples, although the number of examples to retrieve is significantly greater. Indeed, as the training set size $N$ increases, the complexity of the underlying optimization problem increases as well, as the number of possible reconstructions increases exponentially with $N$. However, in the context of DP, considering larger values of $N$ also improves the signal-to-noise ratio. The leaves’ counts also become more informative, as their support grows and allows making more overlaps with those of other trees. Technically speaking, since each modelled training example comes with a set of constraints, the number of such constraints increases accordingly: the search space is larger but also more constrained. This likely explains the better performance of our attack on larger training sets.
@@
TABLE VII: Results of our experiments on the reconstruction attack’s scalability. For varying training set sizes $N$, we report the predictive performances of the built DP RFs, along with the resulting reconstruction error (averaged over 55 different random seeds).
@@
25 0.212 0.023 0.920 0.772
50 0.174 0.010 0.876 0.780
100 0.152 0.013 0.856 0.782
200 0.130 0.009 0.828 0.782
300 0.124 0.006 0.827 0.782
400 0.117 0.006 0.828 0.786
500 0.112 0.006 0.814 0.789
Similar trends can be seen in Figure 10, which displays the anytime performances of our attack. More precisely, we report the reconstruction error value as a function of the solver running time. Note that, at any given time, if some runs (i.e., certain random seed values out of the five considered) have not yet produced a feasible reconstruction, their reconstruction error is replaced by that of the random baseline. This substitution explains the “stairs” effect observed in some plots. As running time increases, the CP solver progressively finds reconstructions with better objective function values (as defined in (9)), which in most cases also correspond to smaller reconstruction errors. This illustrates the alignment between our objective (maximizing the log-likelihood of the inferred DP noise values) and the reconstruction error. The effect of the training set size $N$ is also apparent: reconstructing larger training sets is computationally more challenging, and feasible reconstructions are found more slowly as $N$ increases. Nevertheless, reconstruction errors ultimately tend to converge to better values for larger $N$ due to a more favourable signal-to-noise ratio.
(a) UCI Adult Income dataset
(b) COMPAS dataset
(c) Default of Credit Card Clients dataset
Figure 10: Anytime reconstruction error for the experiments on our attack’s scalability. We report the reconstruction error as a function of the running time, for varying training set sizes NN. For these experiments, all forests are built of $\lvert\mathcal{T}\rvert=10$ trees of depth $5$, with DP budget $\varepsilon=5$. Reconstruction errors are averaged over 5 different random seeds.
Finally, Figure 11 displays the trade-offs between the DP RF training accuracy, test accuracy, and the reconstruction attack performance for various training set sizes. As shown most clearly for the Default of Credit Card Clients dataset in Figure 11(c), increasing the number of training examples $N$ simultaneously (i) decreases the training accuracy, (ii) increases the test accuracy, which indicates improved generalization of the DP RF, and (iii) reduces the reconstruction error achieved by our attack (thereby improving its success). Larger values of $N$ also slightly improve the random baseline, since a larger candidate set makes the alignment step easier and increases the likelihood of partial matches. However, the decrease in reconstruction error obtained by our attack is consistently larger than the decrease observed for the random baseline. This shows that our attack is genuinely exploiting the additional information provided by larger training sets and is not simply benefiting from the alignment artifact. As discussed previously, a likely explanation is that larger training sets lead to higher leaf counts for similar noise magnitudes, which improves the signal-to-noise ratio within the trees. These results illustrate that this stronger signal benefits both the DP RF accuracy and the reconstruction attack.
(a) UCI Adult Income dataset
(b) COMPAS dataset
(c) Default of Credit Card Clients dataset
Figure 11: Average training and test accuracy of $\varepsilon$-DP RFs as a function of the reconstruction error of our attack for varying training set sizes NN. For these experiments, all forests are built of $\lvert\mathcal{T}\rvert=10$ trees of depth $5$, with DP budget $\varepsilon=5$. All values are averaged over 5 different random seeds.
@@ Appendix F Additional Experiments on Partial Reconstruction
We report in Figure 12 the reconstruction error (measured for the unknown attributes) as a function of the number of known attributes. The results for all three datasets consistently show that the knowledge of a number of attributes does not help reconstruct the others. The slight increase in reconstruction error when the number of known attributes increases can be explained by the alignment process: as explained in Section IV-A, to figure out which reconstructed example corresponds to which original one, we first perform a minimum cost matching, before computing the average reconstruction error for each pair of original-reconstructed examples. In this partial reconstruction setup, this matching is still performed on the complete examples’ features vector even though the attack only retrieves part of their attributes. Hence, the improvement in reconstruction error induced by the matching process reduces as the number of unknown attributes decreases.
This observation empirically verifies the theoretical guarantees offered by DP: the protection enforced for the unknown attributes can not be “undone” through postprocessing the DP RF, even with the auxiliary information provided by the known attributes. In other words, in our considered setup, and for our proposed attack, DP prevents so-called linkage attacks [11]. It is worth noting that these results contrast with the ones observed against non-DP RFs, for which knowledge of some of the attributes greatly helped reconstruct the others [9].
@@
Appendix G Additional Experiments on Reconstruction in an Informed Adversary Setting
As detailed throughout the paper, our reconstruction attack is designed to recover the entire training set of a DP-protected RF, with or without prior knowledge of its size $N$. To the best of our knowledge, no existing work has targeted full training-set reconstruction for DP-protected models. As explained in Section V, a related line of work instead considers a simplified informed adversary setting, where the attacker knows N1N-1 training examples and attempts to reconstruct the remaining one by using both the known examples and the model parameters. This setting is interesting because it focuses on reconstructing a single individual, which is closely aligned with the original protection goal of DP, and it provides a meaningful way to compare our attack with existing approaches.
In this appendix, we show that our attack can be adapted to handle this particular setup, and that it outperforms the state-of-the-art.
Adapted CP formulation. Two main modifications to the CP formulation provided within Section III can be done to adapt our attack to the informed adversary setup.
First, since N1N-1 examples are known by the attacker, we simply set their attributes and classes to their true value:
$z_{kc}=1$ $k\in\{1,\dots,N-1\},\quad c\in\mathcal{C},\quad c=c_{k}$ (17)
$z_{kc}=0$ $k\in\{1,\dots,N-1\},\quad c\in\mathcal{C},\quad c\neq c_{k}$ (18)
$\{x_{ki}\}_{i\in\{1,\dots,M\}}=\mathbf{x}_{k}$ $k\in\{1,\dots,N-1\}.$ (19)
More precisely, for each known example $k\in\{1,\dots,N-1\}$, Constraints (17) and (18) fix the one-hot variables $z_{kc}$ to the true class label $c_{k}$, while Constraint (19) fixes its reconstructed attribute vector $\{x_{ki}\}_{i\in\{1,\dots,M\}}$ to the true attribute values $\mathbf{x}_{k}$.
Second, the knowledge of N1N-1 training examples provides additional information about the underlying data distribution, which was not available in the full-dataset reconstruction setting. This information can be leveraged to encourage the reconstructed example to be not only compatible with the DP RF counts and structure, but also consistent with the empirical distribution of the known examples. To incorporate this constraint, we add a regularization term to the objective function that penalizes the average $\ell_{1}$ distance (which corresponds to the Manhattan distance in our binary setting) between the reconstructed example and the known training examples. The resulting objective function is:
$\max\left(-\sum\limits_{k\in\{1,\dots,N-1\}}{\lVert\{x_{Ni}\}_{i\in\{1,\dots,M\}}-\mathbf{x}_{k}\rVert_{1}}\right)+\alpha\left(\sum\limits_{t\in\mathcal{T}}\sum\limits_{v\in\smash{\mathcal{V}_{t}^{L}}}\sum\limits_{c\in\mathcal{C}}\sum\limits_{l=-\gamma}^{\gamma}\log(p_{l})\mathds{1}_{\Delta_{tvc}=l}\right)$ (20)
where the first term acts as a regularizer that favors reconstructions lying close to the support of the empirical data distribution, and the second term is the original log-likelihood objective from (9). From a Bayesian perspective, the regularization term that penalizes the Manhattan distance to the known training examples can be interpreted as introducing a data-dependent log prior on the unknown example. In practice, we rescale the coefficients inside the summations so that both terms have comparable magnitudes, ensuring that their relative importance is controlled solely by the parameter $\alpha$. In our preliminary experiments, we found out that $\alpha=20\varepsilon$ typically leads to good reconstruction results, and we retain this value throughout our experiments. For small privacy budgets $\varepsilon$, this choice places greater emphasis on producing a reconstruction that aligns with the empirical data distribution, whereas larger values of $\varepsilon$ increase the influence of the DP RF structure.
Informed adversary baseline (adapted from [47]). The reconstruction attack studied in [47] considers an informed adversary who knows all training examples except one and has white-box access to the model parameters. Their closed-form attacks for convex models (such as generalized linear models) do not directly transfer to tree ensembles, but their reconstructor-network (RecoNN) strategy for general models can be adapted to DP RFs. Since their original work focuses on image classifiers, and does not consider tabular data or tree-based models, our adaptation to DP RFs required specific design and careful tuning.
In a nutshell, their attack trains a reconstructor network that learns a mapping from model parameters to the unknown training example. To construct training data for the RecoNN, the adversary uses the N1N-1 known training examples (the fixed dataset), together with a set of $N^{\prime}$ shadow examples. For each shadow example, they train a shadow model on the union of the fixed dataset and that shadow example. These shadow models are then serialized into one-dimensional parameter vectors that can be fed to the RecoNN. This procedure yields $N^{\prime}$ pairs of (serialized shadow model, shadow example), which are used to train the RecoNN. Finally, at attack time, the RecoNN receives the serialized parameters of the released model under attack and outputs a prediction for the unknown training example.
Serializing DP RFs into one-dimensional vectors for input to a RecoNN is a key aspect of the adaptation. A natural first approach is to flatten each tree in level-order (that is, using a breadth-first traversal from the root, visiting nodes level by level from left to right). For each internal node, we would store its split attribute and threshold, and for each leaf, its per-class noisy counts. However, this representation performed poorly in practice, even with extensive per-component normalization. The difficulty is that two DP RFs trained on slightly different datasets typically use different split attributes and thresholds at corresponding nodes, so examples traverse the trees differently and land in different leaves. As a result, without specialized architecture, the RecoNN struggled to learn meaningful correspondences between the serialized internal-node parameters and the structure of the underlying trees.
To address this issue, we adopted a simpler serialization scheme. During training, we fix the split attributes and thresholds to those of the observed DP RF, and only serialize the one-dimensional vector of per-class per-leaf noisy counts. This greatly reduces the complexity of the input representation, allowing the RecoNN to focus on learning a mapping from leaf-level signals to the unknown example. We also found that subtracting the contribution of the N1N-1 known examples from these counts improved training stability, and we apply this correction in all experiments. Fixing the tree structure in this way enables the RecoNN to reliably associate positions in the leaf-count vector with the conditions that define each leaf and to learn the corresponding relationship to the unknown training example.
Since the information from the N1N-1 known examples is already embedded in the serialization procedure, we found that training a single RecoNN on a larger auxiliary dataset (disjoint from the training set but drawn from the same distribution) and reusing it to reconstruct each unknown example both improves reconstruction quality and significantly reduces computational cost. In practice, we use the examples that were not selected in the DP RF training set as this auxiliary dataset.
In our experiments, the RecoNN was trained on 1,6001{,}600 pairs of (serialized shadow DP RF, shadow example), using batches of size $32$ for $5$ epochs. The implementation relies on the PyTorch library [65], with a BCEWithLogitsLoss and an AdamW optimizer (learning rate $10^{-3}$, weight decay $10^{-4}$). The RecoNN architecture is a feed-forward neural network with three hidden layers of sizes $1024$, $512$, and $256$. Each hidden layer applies a linear transformation, followed by a ReLU activation and a LayerNorm operation. The final layer is a linear projection that outputs logits for the reconstructed example. To extract the reconstructed example from the RecoNN, we pass the serialized DP RF through the network to obtain a vector of logits. These logits are then transformed into probabilities using a sigmoid activation. Since all attributes are binary, the final reconstructed example is obtained by thresholding each probability at $0.5$.
Simple per-coordinate majority baseline. For very small DP budgets $\varepsilon$ (in particular for $\varepsilon\leq 1$), the noise added to the leaves’ counts dominates the signal, and only limited information can be extracted from them (as was rigorously assessed in Appendix D in the full reconstruction setup). In this regime, both our approach and the baseline from [47] frequently output nearly constant examples that reflect the global data distribution rather than the specific DP RF parameters. To characterize this extreme case more precisely, we introduce a simple baseline that outputs the best constant reconstruction possible with respect to our evaluation metric. More precisely, this baseline selects the example $\mathbf{x}$ that minimizes the average Manhattan distance to all $N$ training examples, which is equivalent to choosing, for each coordinate, the empirical marginal majority among the training set. Since our reconstruction error is also measured using Manhattan distance, this baseline provides a lower bound on the performance of any attack that outputs a constant reconstruction.
Experiments. We focused on a representative DP RF configuration with $\lvert\mathcal{T}\rvert=10$ trees of depth $5$. For each of the three datasets, we varied the DP budget $\varepsilon\in\{0.1,1,5,10,20,30,10^{3}\}$ and ran each configuration with $10$ random seeds. We included the value $\varepsilon=10^{3}$ because, in [47], the attack required very large privacy budgets to succeed. This setting also serves as a proxy for evaluating the attacks in a regime where almost no noise is added to the DP RF parameters.
For each experiment, we trained a DP RF on $N$ training examples, with N=2, [^0] N=2{,}000 (respectively, N=10, [^0] N=10{,}000 and N=20, [^0] N=20{,}000) for the COMPAS (respectively, Default of Credit Card Clients and UCI Adult Income) dataset. Indeed, in this setting, the computational cost of our attack is very small, enabling it to handle DP RFs trained on large datasets. However, at least $N$ examples (disjoint from the training set) must be reserved to form the RecoNN training set required by the implementation of the baseline of [47]. For each run, we randomly sample $100$ training examples and iteratively attempt to reconstruct each of them, using the remaining N1N-1 training examples as known. We imposed a two-hour timeout per experiment (covering the reconstruction of all 100 examples), although in practice all methods completed well within this limit.
Results. The results presented in Figure 13 show that our attack consistently outperforms the baseline from [47]. For all privacy budgets $\varepsilon$, it systematically recovers more information about the missing example, as evidenced by lower reconstruction errors. This can be explained by the fact that our attack is tailored to exploit the full structure of a DP RF, whereas the baseline is model-agnostic (even though we made a significant effort to design a serialization and RecoNN training procedure that is effective in this setting).
(a) UCI Adult Income dataset
(b) COMPAS dataset
(c) Default of Credit Card Clients dataset
Figure 13: Results of our reconstruction experiments in the informed adversary setup. We report the reconstruction error for the single unknown example as a function of the DP budget $\varepsilon$ used to train the DP RFs. In these experiments, all forests contain $\lvert\mathcal{T}\rvert=10$ trees of depth $5$. The reconstruction attacks are run iteratively over a random subset of the training examples, and the reported reconstruction errors are averaged across these examples and across $10$ random seeds. Standard deviations are also shown.
We observe that for privacy budgets $\varepsilon>5$ (or $\varepsilon\geq 5$ for the UCI Adult Income and Default of Credit Card Client datasets), our attack is able to retrieve meaningful information about the unknown training example. For smaller budgets ($\varepsilon\leq 1$ for our attack, or $\varepsilon\leq 10$ for the RecoNN baseline), only distributional information is recovered, which is consistent with our observations in the full-dataset reconstruction setting (Table II), as indicated by the fact that the reconstruction error is no better than that of the simple per-coordinate majority baseline. These results highlight that for moderate ($\varepsilon\in\{5,10\}$) to large ($\varepsilon\geq 20$) DP budgets, our approach is able to recover meaningful information by thoroughly exploiting the DP RF parameters. The computational cost of our attack is also very low in this setting, with an average runtime of approximately 4 seconds (respectively, 9 and 13 seconds) for the COMPAS (respectively, Default of Credit Card Clients and UCI Adult Income) dataset.
(a) UCI Adult Income dataset
(b) COMPAS dataset
(c) Default of Credit Card Clients dataset
Figure 14: Results of our reconstruction experiments in the informed adversary setup. Compared to Figure 13, results are reported only for our informed adversary, with the average reconstruction error (and standard deviation) being computed and reported separately for inliers and outliers.
(a) UCI Adult Income dataset
(b) COMPAS dataset
(c) Default of Credit Card Clients dataset
Figure 15: Results of our reconstruction experiments in the informed adversary setup. Compared to Figure 13, results are reported only for the baseline informed adversary [47], with the average reconstruction error (and standard deviation) being computed and reported separately for inliers and outliers.
(a) UCI Adult Income dataset
(b) COMPAS dataset
(c) Default of Credit Card Clients dataset
Figure 16: Results of our reconstruction experiments in the informed adversary setup. We report the average number of perfectly reconstructed examples as a function of the DP budget, for the our proposed informed adversary [47], considering either inlier or outlier unknown examples.
(a) UCI Adult Income dataset
(b) COMPAS dataset
(c) Default of Credit Card Clients dataset
Figure 17: Results of our reconstruction experiments in the informed adversary setup. We report the average number of perfectly reconstructed examples as a function of the DP budget, for the baseline informed adversary [47], considering either inlier or outlier unknown examples.
Figures 14 and 15 refine these results by reporting them separately for inlier and outlier examples, for our informed adversary and the informed adversary baseline adapted from [47]. As detailed in Section IV-B, each training example was classified as either inlier or outlier using isolation forests as implemented in the scikit-learn [23] Python library, with their default hyperparameters.
The observed trends are consistent with Result 5 (and Table III) from Section IV-B in the full reconstruction setup. Indeed, even in the informed adversary setting, the average reconstruction error is consistently lower for inliers. This can be explained by the fact that the distributional information extracted by the attack facilitates the reconstruction of inliers, while offering only limited benefit for outliers. The same qualitative behavior is observed for the baseline informed adversary.
These results further confirm our previous observations. For privacy budgets $\varepsilon>5$ (or $\varepsilon\geq 5$ for the Default of Credit Card Clients dataset), our attack can recover meaningful information about the unknown training example, even when it lies outside the data distribution. This indicates that, in these privacy regimes, the attack infers example-specific information beyond distributional patterns.
The superiority of our adversary is also apparent, as the baseline is only able to extract non-trivial information about outliers for substantially larger privacy budgets ($\varepsilon\geq 20$). The reconstruction error of outliers also quickly converges towards that of inliers (close to 0) (for $\varepsilon=20$ or $\varepsilon=30$, depending on the dataset) with our proposed attack. With the considered baseline, this convergence is considerably slower (as even with $\varepsilon=1000$, a small gap remains).
For smaller privacy budgets ($\varepsilon<5$), the inferred information remains purely distributional and therefore only helps the reconstruction of inliers for both adversaries, as expected.
Figures 16 and 17 confirm these trends by reporting the proportions of inlier and outlier examples that are perfectly reconstructed for different DP budgets, for our informed adversary and the informed adversary baseline adapted from [47]. In particular, our informed adversary is often able to perfectly reconstruct the unknown example even when it lies outside the data distribution and when the DP budget is moderate, and the proportion of perfectly reconstructed unknown examples increases rapidly with the DP budget $\varepsilon$.
Altogether, these results show that tight DP budgets are required to prevent DP RFs from leaking dataset-specific patterns, including information about out-of-distribution examples, which is consistent with our findings in the full reconstruction setup presented in the main paper. They further indicate that the DP budget $\varepsilon$ required to prevent successful reconstruction in the informed adversary setting depends on both the dataset and the attack considered. This motivates the use of our informed adversary as a meaningful vulnerability audit, as it models all the information provided through the DP RF structure and parameters, and empirically outperforms the baseline attack.